# Enhance Image-to-Image Generation with LLaVA Prompt and Negative Prompt

### v1

@@ Abstract
The field of artificial intelligence is experiencing a surge in the development of multifaceted assistants capable of solving many real-life problems. Large language models play an important part of it, showing their great in multiple fields such as natural language processing [1, 2, 3, 4, 5, 6, 7] and computer vision [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Among those fields, one such example is Stable Diffusion [18], a powerful model that generates novel and imaginative images based on user-provided starting points. However, relying solely on the input image presents limitations. The generated image may deviate significantly from the user’s intent, resulting in outputs that lack the desired level of control or fidelity. In addition, the inaccuracy and instability of generated images from large language models are also problems observed in past studies that necessitate careful consideration and mitigation strategies moving forward [19].
@@ II-B Image-to-Image Generation with Prompts
Our methodology utilizes the StableDiffusionImg2ImgPipeline, a component within Stable Diffusion 2.0. This pipeline employs a diffusion-denoising mechanism, granting us fine-grained control over the image editing process. We incorporate input images alongside prompts generated by LLaVA. These prompts aim to capture the essence of the input image. Additionally, negative prompts are employed to guide the model away from unintended visual elements. By combining the input image, LLaVA-generated prompts, and negative prompts, we instruct the image-to-image pipeline to generate a new image that closely resembles the original input image but incorporates the specified modifications [28, 29, 30, 31]. The fine-grained control offered by text prompts and the integration with LLaVA-generated prompts provide a robust approach for manipulating and generating new visuals [32, 33, 34, 35, 36, 37, 38, 39, 40].
@@ III-D Comparison w/wo Prompt
Figure 2 (b) and (c) visually demonstrate the distinction between images generated with and without LLaVA prompts. To quantify this observed difference in image similarity, we calculate the image similarity metrics, including Root Mean Square Error (RMSE), Peak Signal-to-Noise Ratio (PSNR), Feature-based Similarity Index (FSIM), Structural Similarity Index (SSIM), Universal Image Quality Index (UIQ), and Signal to Reconstruction Error Ratio (SRE) [41, 42, 43, 44]. The results, presented in Table 1, substantiate the qualitative observations, indicating that LLaVA-generated prompts contribute to the generation of more similar images.
@@ IV CONCLUSIONS AND FUTURE WORK
For the future work, we envision further exploration into fine-tuning LLaVA prompts for increased control over the creative process. By providing more detailed specifications within the prompts, we aim to guide the image-to-image pipeline towards emphasizing specific aspects of the input image while maintaining creative freedom in other areas [45, 46, 47, 48, 49]. This refined approach holds promise for achieving a delicate balance between similarity and creativity.

### v3

# Enhance Image-to-Image Generation with LLaVA-generated Prompts

@@ Abstract
The field of artificial intelligence is experiencing a surge in the development of multifaceted assistants capable of solving many real-life problems. Large language models play an important part of it, showing their great in multiple fields such as natural language processing [1, 2, 3] and computer vision [4, 5, 6, 7, 8, 9, 10, 11, 12]. Among those fields, one such example is Stable Diffusion [13], a powerful model that generates novel and imaginative images based on user-provided starting points. However, relying solely on the input image presents limitations. The generated image may deviate significantly from the user’s intent, resulting in outputs that lack the desired level of control or fidelity. In addition, the inaccuracy and instability of generated images from large language models are also problems observed in past studies that necessitate careful consideration and mitigation strategies moving forward [14].
@@ II-B Image-to-Image Generation with Prompts
Our methodology utilizes the StableDiffusionImg2ImgPipeline, a component within Stable Diffusion 2.0. This pipeline employs a diffusion-denoising mechanism, granting us fine-grained control over the image editing process. We incorporate input images alongside prompts generated by LLaVA. These prompts aim to capture the essence of the input image. Additionally, negative prompts are employed to guide the model away from unintended visual elements. By combining the input image, LLaVA-generated prompts, and negative prompts, we instruct the image-to-image pipeline to generate a new image that closely resembles the original input image but incorporates the specified modifications [23, 24]. The fine-grained control offered by text prompts and the integration with LLaVA-generated prompts provide a robust approach for manipulating and generating new visuals [25, 26, 27, 28].
@@ III-D Comparison w/wo Prompt
Figure 2 (b) and (c) visually demonstrate the distinction between images generated with and without LLaVA prompts. To quantify this observed difference in image similarity, we calculate the image similarity metrics, including Root Mean Square Error (RMSE), Peak Signal-to-Noise Ratio (PSNR), Feature-based Similarity Index (FSIM), Structural Similarity Index (SSIM), Universal Image Quality Index (UIQ), and Signal to Reconstruction Error Ratio (SRE) [29]. The results, presented in Table 1, substantiate the qualitative observations, indicating that LLaVA-generated prompts contribute to the generation of more similar images.
@@ IV CONCLUSIONS AND FUTURE WORK
For the future work, we envision further exploration into fine-tuning LLaVA prompts for increased control over the creative process. By providing more detailed specifications within the prompts, we aim to guide the image-to-image pipeline towards emphasizing specific aspects of the input image while maintaining creative freedom in other areas [30, 31]. This refined approach holds promise for achieving a delicate balance between similarity and creativity.